On January 1, 2018, Ms. Fawn Halpern owned 9500 shares of Zunit Inc., a Canadian public company. These shares were acquired several years ago at a cost of $23.50 per share, a total investment of $223,250.

During 2018, the following transactions occur:

  • On February 1, she gives 3500 of these shares to her common-law partner Melvin Young. On this date, the shares are trading at $25.00 per share. Fawn does not elect out of ITA 73(1).
  • On March 1, she gives the remaining 6000 shares to her 15 year old daughter Clare. At this time the price of the shares has fallen to $22.50.
  • On July 1 and September 1, the Zunit Inc. shares pay an eligible dividend of $0.80 per share.
  • On December 1, both Melvin and Clare sell all of the shares they have received for $26.25 per share.

Required:

A. Determine the total Net Income for Tax Purposes to be recorded for the 2018 taxation year by Fawn, her common-law partner Melvin, and her daughter Clare, as a result of the preceding transactions.

B. How would your answer change if Fawn died on July 15, 2018?
Answer #1

Number of shares owned by Ms. Fawn of Zunit Inc is equal to 9500 shares

cost price per share = $23.50 Total investment of $ 223250 (9500* 23.50)

During the year she gave 3500 shares to her common -law partner Melvin @25 per share and rest 6000 shares to her daughter Clare at $22.50

Total Net Income for tax purpose for Ms. Fawn will be Investment amount from sale of shares subtracting the initial investment value.

Proceeds from sale of shares to Melvin and Clare = (3500* 25) + (6000* 22.50) i.e. 87500+135000 equals to 222,500

Therefore Ms. Fawn has incurred Net loss of $750 (223250-222500) from the sale of 9500 to Melvin and Clare respectively i.e. 3500 and 6000 shares.

Calculation of Total net Income for Tax purposes for Melvin

Total initial Investment = 3500* 25 = $87500

On July 1 and September 1, company declared $0.80 dividend per share

Dividend payment received by Melvin = (3500* 0.8) *2 times i.e. $5600

Sale of Shares on 1st December = 3500* 26.25 = 92750

Net Income for Melvin = Sale value of shares + dividend received - Initial investment i.e. 92750+5600 - 87500 = $10850

Calculation of Total Net Income for Clare

Total Initial investments = 6000* 22.50 which is equal to $135000

Dividends received = (6000* 0.8) * 2 = $9600

Sale of shares on 1st December = 6000* 26.25 which is equal to $157500

Net Income for Clare = Sales value of 6000 shares + dividend received - Initial Investment i.e. 157500+ 9600 - 135000 = $32100

B) No answer will remain the same in case Ms. Fawn died on 15 July 2018 as she has already sold her entire 9500 shares to Melvin and Clare during the year and dividend payment of $0.8 per share twice in a year will be received by Melvin and Clare in proportion of their share holding. Hence there will no change in the answer if Ms. Fawn died on 15 July. Rights of shares has already being transferred to Melvin and Clare and now they are owner of 3500 and 6000 shares as on 15 July. Ms. Fawn has incurred a net loss of $750 from the sale of her 9000 shares to Melvin and Clare. Also she is not entitled to any dividend payments during the year as shares has already been transferred to Melvin and Clare before declaration of dividends.

Later both of them have sold their shares at the market value of $26.25 per share